OpenStudy (anonymous):

(multiple choice) can anyone help me identify the real roots and state their multiplicity x^4+2x^23x^2-24x+144=0? ( :

OpenStudy (anonymous):

root -4 multipicity 2, root 3multiplicity 2 root -4 multipicity 1, root 3multiplicity 2 root -4 multiplicity 2, root 3multiplicity 1 root -4 multipicity 1, root3 multiplicity 1

OpenStudy (anonymous):

you gotta be kidding? a polynomial of degree 4??

OpenStudy (anonymous):

i guess you are told that \(-4\) is a root so you can factor out the \(-4\) as \(x+4\) and see what is left then try it again me, i would cheat

OpenStudy (anonymous):

I can't see the numbers you typed. It shows "{math processing Error}"

OpenStudy (anonymous):

before we try it, there is a typo in your question the thing you wrote above \(x^4+2x^3-2x^2-24x+144\) has no zeros

OpenStudy (anonymous):

refresh your browser, it is not reading the latex correctly

OpenStudy (anonymous):

check your question again that sucker has no zeros

OpenStudy (anonymous):

if I plug in 4 i get 464 as the sum. The zers is after the equal sign

OpenStudy (anonymous):

I don't understand the multiplicity part :o

OpenStudy (anonymous):

yes, there is a typo in the question

OpenStudy (anonymous):

hm, It's straight from an assignment I'm working on. I'm not sure what the typo is

OpenStudy (anonymous):

please check it again the thing you wrote has no real zeros once we get the right thing i can explain the "multiplicity" part, but not until we get the right expression to find the zeros of

OpenStudy (anonymous):

if that is exactly what is written, then there is a mistake in the question so you cannot do it

OpenStudy (anonymous):

woops I wrote 3-2x^2 it is 23x^2

OpenStudy (anonymous):

x^4+2x^3-23x^2-24x+144=0

OpenStudy (anonymous):

whew now we can do it!

OpenStudy (anonymous):

haha I appreciate your patience ( :

OpenStudy (anonymous):

actually we have to do very little but before we do it, would you like the answer?

OpenStudy (anonymous):

yes! :D

OpenStudy (anonymous):

you see where it says \((x-3)^2 (x+4)^2 = 0\) ?

OpenStudy (anonymous):

yes those are the factors of 12

OpenStudy (anonymous):

that tells you the zeros are \(3\) and \(-4\) and the fact that both of those factors has a exponent of \(2\) means the "multiplicity" of those zeros is 2

OpenStudy (anonymous):

by "factor" i don't mean "factor of the numbers" i mean the factors are \(x-3\) and \(x+4\) the exponent is the multiplicity that shows up in the graph in the fact that the curve touches the \(x\) axis at those zeros but does not cross the axis

OpenStudy (anonymous):

thank you : ) I send love into the universe for ya <3

OpenStudy (anonymous):

lol thanks i didn't show you how to do it without cheating though we can still do that if you like

OpenStudy (anonymous):

yeah ( :

OpenStudy (anonymous):

ok first off it is almost impossible to find the zeros of a 4th degree polynomial but in this case you are given possible answers, and each answer has a 3 in it so you can safely assume a zero is 3 and factor it out . i.e. factor \[x^4+2x^3-23x^2-24x+144\] as \[(x-2)(something)\] and the easiest way to do that is by synthetic division

OpenStudy (anonymous):

damn typo, i meant \[x^4+2x^3-23x^2-24x+144\]\[(x-3)(something)\]

OpenStudy (anonymous):

if you know how to do synthetic division this is quite easy do you know it?

OpenStudy (anonymous):

yes. bring down 1 * 3 3+2=5 5*3= 15 -23+15= -8 -8*3= -24 -24+-24 = -48??? I think I did something wrong

OpenStudy (anonymous):

woops I forgot to multiply -48 by 3 and add it to 144 which leave a remainder of zero

OpenStudy (anonymous):

yes you should have a remainder of zero because it factors, i.e. because if \(3\) is a zero, then you know \(x-3\) is a factor

OpenStudy (anonymous):

what is your final answer?

OpenStudy (anonymous):

(x-3)(x+4) x=3 x=-4?

OpenStudy (anonymous):

oh no i mean after you factor you get \[(x-3)(something)\] what was the "something"?

OpenStudy (anonymous):

x+4?

OpenStudy (anonymous):

after you did the synthetic division, what did you get?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

2, 3, -5, 0

OpenStudy (anonymous):

lol, not as the remainder! as the other part of the factor it should have been \(x^3+5 x^2-8 x-48\)making \[(x-3)(x^3+5 x^2-8 x-48)\]in factored form

OpenStudy (anonymous):

in any case, you have to repeat this process with \(3\) when you do it again, the last factor will be \(x^2+8x+16\) which is \((x+4)^2\) making the whole thing factor as \((x-3)^2(x+4)^2\) as we saw originally, and you are done

OpenStudy (anonymous):

thank you ( : I understand it. I'm just confusing different problems :o

OpenStudy (anonymous):

oh i see that you got it when you did the synthetic division you got the coefficients as \(1,5,-8,-48\) making it what i wrote above it is long and annoying, that is why i like to cheat

OpenStudy (anonymous):

yw